OpenStudy (anonymous):

Use basic identities to simplify the expression. sin^2θ + tan^2θ + cos^2θ sec^2θ cos^3θ sin θ tan^2θ

OpenStudy (anonymous):

u know the double angle equations?

OpenStudy (anonymous):

wait, did u mean \[\sin ^{2}\theta + \tan ^{2} \theta + \cos^{3}\theta?\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

wait no that's wrong

OpenStudy (anonymous):

cos 2 thetha sorry, my typo...

OpenStudy (anonymous):

oh yeah then that's right again :O sorry for the confusion!

OpenStudy (anonymous):

sin^2u + cos^2u = 1 and 1 + tan^2u = sec^2u

OpenStudy (anonymous):

\[using \cos 2x = \cos^{x} - \sin^{x},\] \[\sin^{2}x = \cos^{2}x - \cos 2x\] \[\tan ^{2}x = \frac{ \sin^{2}x }{ \cos^{2}x }\]

OpenStudy (anonymous):

oooh okay I got tan^2θ the first time I tried it and I wasn't sure if it was right! Thanks!